# Detecting Toxic Flow

### v1

@@ 3.2 Deploy phase: estimating last-layer parameters and feature-transform parameters
where ${\boldsymbol{\mu}}_{t}$, ${\boldsymbol{\Gamma}}_{t}$ are the estimated mean and covariance of the projected-hidden-layer parameters at time $t$ and ${\boldsymbol{\nu}}_{t}$, ${\boldsymbol{\Sigma}}_{t}$ are the estimated mean and covariance matrix of the last-layer parameters.
See A for a proof.
4 Model evaluation
We employ the methodology developed in the previous section with the following configuration. The NNet for PULSE is an MLP with three hidden layers, 100 units in each layer, and ReLU activation function. The number of epochs $E$ is 850, we skip the first $n=50$ iterations of the optimisation procedure, the subspace dimension is $d=20$, the learning rate is $\alpha=10^{-7}$, and we store gradients every $k=4$ steps. Under this configuration, the MLP has 38,70038,70038 , 700 units which we estimate during the warmup stage. For the deployment stage, PULSE updates 120120120 degrees of freedom; this accounts for less than half a percent of all parameters updated during the warmup stage. From a practical perspective, a single step of the algorithm (during the deployment stage) incurs in a memory cost of $O((L+D)^{2})$. In this paper, an update requires less than 1mb of memory because each unit is a 32bit float. Conversely, if we were not to employ the subspace approach a single step would require 190gb of memory making it infeasible to run on typical GPU devices.
4.1 Benchmarks
We compare the performance of four methods: PULSE, logistic regression (LogR), random forests (RF), and a recursively updated maximum-likelihood estimator of a Bernoulli-distributed random variable (MLE). With logistic regression, the probability that a trade is toxic is
where ${\bf w}_{0}$ is estimated maximising the log-likelihood using L-BFGS; see Liu and Nocedal, (1989).
Next, RF is a bootstrap-aggregated collection of de-correlated trees. A prediction if a trade is toxic is done by averaging over the trees, see Section 15.1 of Hastie et al., (2001).
Further, we model the unconditional probability of a toxic trade as a Bernoulli-distributed random variable whose mass function is given by
$p(y\,|\,\pi)=\text{Bern}\Big{(}y\,|\,\pi\Big{)},\qquad y\in\{0,1\}\,.$ (4.2)
The maximum likelihood estimator of the parameter $\pi$, given a collection $\{z_{1},\ldots,z_{N}\}$ of Bernoulli-distributed samples, where $z_{n}\in\{0,1\}$ is given by
$\pi_{\text{MLE}}=\frac{1}{N}\,\sum_{n=1}^{N}\mathbbm{1}(z_{n}=1)\,;$
We consider two sets of models: one for buy orders and another for sell orders trained on all clients; in 7.3, as part of the robustness analysis, we study the performance when we train one model per client ID.
5 Deployment of online methods
@@
Next, we discuss how we evaluate the performance of the two online methods (MLE and PULSE) with asynchronous data. In classical filtering problems, as soon as new information arrives at, say, time $t_{i}$, the parameters $\bm{\theta}_{t_{i}}$ of the model are updated. Next, when a new trade arrives at time $t_{i+1}$, one uses the parameters $\bm{\theta}_{t_{i}}$ to estimate if a trade will be toxic. In our setting, however, an update at time $t_{i+1}$ with $\bm{\theta}_{t_{i}}$ is only possible if $t_{i+1}>t_{i}+\mathfrak{G}$. Otherwise, we use $\bm{\theta}_{t_{j}}$ to predict the probability of a toxic trade, with $j=\operatorname*{arg\,max}_{k}t_{i+1}>t_{k}+\mathfrak{G}$. Figure 5 illustrates this procedure.
@@
Figure 5: Asynchronous predict-update steps: trades arrive at irregular times $\{t_{i}\}_{i}$. An update to the model is only possible if $t_{i+1}>t_{i}+\mathfrak{G}$, when we know if the trade was toxic. In this example, the model parameters $\bm{\theta}_{0}$ are known at time $t_{1}$, when a new trade arrives. When a second trade arrives, at time $t_{2}$, we do not know if the previous trade was toxic at $t_{1}$, so we use the model weights $\bm{\theta}_{0}$ to make a prediction. The next trade arrives at time $t_{3}>t_{1}+\mathfrak{G}$, so we use $\bm{\theta}_{1}$ to make a prediction. Finally, multiple trades arrive consecutively at times $t_{4}$, $t_{5}$, and $t_{6}$, in which case we use $\bm{\theta}_{2}$. The last trade to arrive at time $t_{7}$ uses $\bm{\theta}_{6}$. In this example $\bm{\theta}_{3}$, $\bm{\theta}_{4}$, and $\bm{\theta}_{5}$ were never used to make a prediction because there are no trades between $t_{3}$ and $t_{6}$.
We employ the asynchronous online updating for PULSE  and MLE. We select hyperparameters over the warmup stage. Figure 6 shows how PULSE updates model parameters based on: current model parameters $\bm{\theta}$, features $\bf x$, and outcome $y$. Here, $\bm{\theta}=({\boldsymbol{\psi}},{\bf w})$ are the model parameters one uses to produce $p(y=1\,|\,{\bf x},{\bm{\theta}})$, with which we compute the prediction $\hat{y}$. We employ the predictions $\hat{y}$ and the outcomes $y$ to compute the accuracy defined above.
$s$$s+\mathfrak{G}$$t$${\bf x}_{s}$$\boldsymbol{\theta}_{s}$$\hat{y}_{s}$$\boldsymbol{\theta}_{s+\mathfrak{G}}$$y_{s}$${\bf x}_{t}$$\boldsymbol{\theta}_{t}$$\hat{y}_{t}$
Figure 6: PULSE update procedure. For simplicity, we take $s+\mathfrak{G}<t$.
@@ 5.1 Model for decision making
$\delta^{\pm*}=\operatorname*{arg\,max}_{\delta^{\pm}\{0,1\}}\mathbb{E}\left[\underbrace{\overbrace{\pm\delta^{\pm}\,(S\pm\mathfrak{S})}^{\text{cash flow}}+\overbrace{(S\pm\eta\,Z)\left(Q\mp\delta^{\pm}\right)}^{\text{inventory valuation}}}_{\text{mark-to-market}}\underbrace{-\phi\,\left(Q\mp\delta^{\pm}\right)^{2}}_{\text{inventory penalty}}\right]\,,$ (5.1)
where $Z$ is a Bernoulli random variable with parameter $p^{\pm,\text{M}}$, and $\phi\geq 0$ is an inventory penalty parameter. Intuitively, the broker optimises the expected value of her mark-to-market adjusted by a quadratic penalty on inventory. The solutions to (5.1) are
$\delta^{\pm*}=\mathds{1}\left(\frac{\mathfrak{S}}{\eta}-\frac{\phi}{\eta}\pm\frac{2\,\phi}{\eta}\,Q>p^{\pm,\text{M}}\right)=\mathds{1}\left(\mathfrak{p}\pm\Phi\,Q>p^{\pm,\text{M}}\right)\,,$ (5.2)
where $\mathfrak{p}:={\mathfrak{S}}/{\eta}-\phi/{\eta}$ and $\Phi:=2\,\phi/\eta$. We call $\Phi$ the inventory aversion parameter and we call $\mathfrak{p}$ the cutoff probability. The strategy internalises trades when the prediction $p^{\pm,\text{M}}$ is smaller than the cutoff probability $\mathfrak{p}$ adjusted by the inventory of the broker $Q$ and the inventory aversion parameter $\Phi$. When either
$\mathfrak{p}+\Phi\,Q=p^{+,\text{M}}\qquad\text{ or }\qquad\mathfrak{p}-\Phi\,Q=p^{-,\text{M}}\,,$
the broker is indifferent between internalising or externalising the trade in the market; this happens with probability zero. Below, we study the case when $\Phi=0$ in more detail. In 7.1 we explore the case when $\Phi>0$.
@@ 5.2 Internalise-externalise strategy
Motivated by the mathematical framework in Subsection 5.1, below we introduce a family of predictions of toxicity based on the cutoff probability $\mathfrak{p}\in[0,1]$.
@@
We are interested in the predictive performance of the models as we vary $\mathfrak{p}$. To this end, let $y_{t_{n}}\in\{0,1\}$ denote if a trade executed at time $t_{n}$ was toxic at time $t_{n}+\mathfrak{G}$ ($y_{t_{n}}=1$ if toxic and $y_{t_{n}}=0$ otherwise). We employ the true positive rate and the false positive rate, which we define below.
@@
The area under an ROC curve, called AUC, is used in the machine learning literature to compare classifiers; see e.g., Fawcett, (2006). Intuitively, the AUC is a measure to quantify a classifier’s ability to distinguish between toxic and benign trades.
@@ 5.3 Model comparison
In this section, we analyse the AUC of the methods for varying toxicity horizons. Figure 8 shows the daily density estimate of AUC for each of the methods. That is, for each method, we compute the AUC for the sequence of trades of each day and show a density plot of such values; recall that the deploy window is between 1 August 2022 and 21 October 2022.
Figure 8: Left panel shows the kernel density estimate plot of daily AUC by toxic horizon. Right panel shows the daily AUC of all models for a toxicity horizon of 30s. The dashed vertical lines correspond to the mean daily value. EUR/USD currency pair over the period 1 August 2022 to 21 October 2022.
PULSE has the highest average AUC among the four methods for all toxicity horizons. Furthermore, as the toxic horizon increases, the AUC of PULSE and RF decreases, but the outperformance of PULSE over RF increases. MLE and LogR have the poorest performance.
Figure 9 shows the five-day exponentially-weighted average of the AUC for each day. We see that, whereas RF, PULSE, and MLE obtain their maximum values at the beginning of the deploy period and then decay over time, PULSE, on the other hand, maintains a steady performance. This is because PULSE updates its parameters with each new observation.
@@
Figure 10 shows a five-day exponentially-weighted moving average (with decay 1/31/31 / 3) of the AUC for the various toxicity horizons across time.
@@
We see that as the toxicity horizon increases, the maximum AUC decreases. In all cases, the AUC for PULSE remains stable througout the deploy period.
6 Trade Prediction Effectiveness and Missed Opportunities
We employ data for the EUR/USD currency pair over the period 1 August 2022 to 21 October 2022 to test the internalisation-externalisation strategy. As above, we ignore inventory aversion (i.e., $\Phi=0$); in Section 7 we perform robustness checks for the inventory aversion and study how the performance of the internalisation-externalisation strategy changes as the broker’s inventory aversion increases.
Figure 11 reports the PnL ($y$-axis) of the internalised trades and the avoided PnL ($x$-axis) of the externalised trades when the broker uses the internalisation-externalisation strategy (5.2) with cutoff probability $\mathfrak{p}\in\{0.05,0.15,0.25,\ldots,0.95\}$ and for PULSE, MLE, LogR, and RF. The broker starts with zero inventory at the beginning of 1 August 2022 and she crosses the spread to unwind the internalised trades at the end of the toxicity horizon, to later report the avoided PnL. We assume there is no cost incurred in externalising trades and we keep track of the PnL they would have had over the toxicity horizon. The inventory is in euros (€) and the PnL in dollars (\$). We take each trade to be for the median quantity which is €2,000 in our dataset. When a trade is toxic, the median cost to unwind the position is \$0.00007-\\$0.00007- \$ 0.00007 and when the trade is not toxic, the median profit is \$0.0008\\$0.0008\$ 0.0008.
Figure 11: PnL and avoided loss for various toxicity horizons and for $\mathfrak{p}\in\{0.05,0.15,0.25,\ldots,0.95\}$.
For each of the toxicity horizons, the internalisation-externalisation strategy informed by PULSE attains the highest PnL and the lowest avoided loss. This is indicated with the red dots joined by the red dash line. The best possible outcome is for the ten-second toxicity horizon. Next, Figure 12 shows how the percentage of internalised volume depends on the cutoff probability.

### v2

@@ 4.1.2 Deploy phase: online estimation of last-layer and subspace-feature-transform parameters
where ${\boldsymbol{\mu}}_{n}$, ${\boldsymbol{\Gamma}}_{n}$ are the estimated mean and covariance of the projected-hidden-layer parameters at step $n$, and ${\boldsymbol{\nu}}_{n}$, ${\boldsymbol{\Sigma}}_{n}$ are the estimated mean and covariance matrix of the last-layer parameters.
As a corollary, a variant of PULSE can be derived to model any other member of the exponential family by replacing the mean and covariance of the target distribution of choice. See B for a proof of Theorem 2.
5 Asynchronous learning and decision making
@@
Next, we discuss how we deploy and evaluate the performance of the online PULSE methodology with asynchronous data. In classical filtering problems, as soon as new information arrives at, say, time $t_{i}$, the parameters $\bm{\theta}_{t_{i}}$ of the model are updated. Next, when a new trade arrives at time $t_{i+1}$, one uses the parameters $\bm{\theta}_{t_{i}}$ to estimate if a trade will be toxic. In our setting, however, an update at time $t_{i+1}$ with $\bm{\theta}_{t_{i}}$ is only possible if $t_{i+1}$ is greater than the time of last trade $t_{i}$ plus the toxicity horizon $\mathcal{G}>0$, i.e., $t_{i+1}>t_{i}+\mathcal{G}$. Otherwise, we use $\bm{\theta}_{t_{j}}$ to predict the probability of a toxic trade, with $j=\operatorname*{arg\,max}_{k}t_{i+1}>t_{k}+\mathcal{G}$. Figure 6 illustrates this procedure.
@@
Figure 6: Asynchronous predict-update steps: trades arrive at irregular times $\{t_{i}\}_{i}$. An update to the model is only possible if $t_{i+1}>t_{i}+\mathcal{G}$, when we know whether the trade was toxic or benign. In this example, the model parameters $\bm{\theta}_{0}$ are known at time $t_{1}$, when a new trade arrives. When a second trade arrives, at time $t_{2}$, we do not know whether the previous trade was toxic or benign at $t_{1}$, so we use the model weights $\bm{\theta}_{0}$ to make a prediction. The next trade arrives at time $t_{3}>t_{1}+\mathcal{G}$, so we use $\bm{\theta}_{1}$ to make a prediction. Finally, multiple trades arrive consecutively at times $t_{4}$, $t_{5}$, and $t_{6}$, in which case we use $\bm{\theta}_{2}$. The last trade to arrive at time $t_{7}$ uses $\bm{\theta}_{6}$. In this example $\bm{\theta}_{3}$, $\bm{\theta}_{4}$, and $\bm{\theta}_{5}$ were never used to make a prediction because trades did not arrive during the period $t_{3}+\mathcal{G},t_{6}+\mathcal{G})$.
We employ the asynchronous online updating for PULSE and MLE and select hyperparameters over the warmup stage. Figure 7 shows how PULSE updates model parameters based on: current model parameters $\bm{\theta}$, features $\bf x$, and outcome $y$. Here, $\bm{\theta}=({\boldsymbol{\psi}},{\bf w})$ are the model parameters one uses to produce $p(y=1\,|\,{\bf x},{\bm{\theta}})$, with which we compute the prediction $\hat{y}$. We employ the predictions $\hat{y}$ and the outcomes $y$ to compute the accuracy defined above.
[…]
Figure 7: PULSE update procedure. For simplicity, we take $s+\mathcal{G}<t$.
@@ 5.1 Model for decision making
For method $\text{M}\in\{\text{PULSE},\text{ LogR},\text{ RF},\text{ MLE}\}$, let $p^{+,\text{M}}\in(0,1)$, denote the probability that a buy order will be toxic and let $p^{-,\text{M}}$ denote the probability that a sell order will be toxic. Note that $p^{+,\text{M}}+p^{-,\text{M}}$ does not necessarily add to $1$. Let $\mathcal{S}/2>0$ denote the half bid-ask spread and let $\eta>0$ denote the shock to the midprice $S$ if the trade is toxic; here we assume that $\eta\in(\mathcal{S},\infty)$. The broker controls $\delta^{\pm}\in\{0,1\}$. When $\delta^{\pm}=0$ the broker externalises the trade and when $\delta^{\pm}=1$ the broker internalises the trade. The inventory of the broker is $Q\in\mathbb{R}$; when $Q>0$ the broker is long and when $Q<0$ the broker is short. Assume all trades are for one unit of the asset. Then, the broker solves
$\delta^{\pm*}=\operatorname*{arg\,max}_{\delta^{\pm}\{0,1\}}\mathbb{E}\left[\underbrace{\overbrace{\pm\delta^{\pm}\,(S\pm\mathcal{S}/2)}^{\text{cash flow}}+\overbrace{(S\pm\eta\,Z)\left(Q\mp\delta^{\pm}\right)}^{\text{inventory valuation}}}_{\text{mark-to-market}}\underbrace{-\phi\,\left(Q\mp\delta^{\pm}\right)^{2}}_{\text{inventory penalty}}\right]\,$, (5.1)
where $Z$ is a Bernoulli random variable with parameter $p^{\pm,\text{M}}$, and $\phi\geq 0$ is an inventory penalty parameter. Intuitively, the broker optimises the expected mark-to-market value of her portfolio after internalising the trade adjusted by a quadratic penalty on inventory. The solutions to (5.1) are
$\delta^{\pm*}=\mathds{1}\left(\frac{\mathcal{S}/2}{\eta}-\frac{\phi}{\eta}\pm\frac{2\,\phi}{\eta}\,Q>p^{\pm,\text{M}}\right)=\mathds{1}\left(\mathfrak{p}\pm\Phi\,Q>p^{\pm,\text{M}}\right)\,$, (5.2)
where $\mathfrak{p}:={\mathcal{S}}/{2\,\eta}-\phi/{\eta}$ and $\Phi:=2\,\phi/\eta$. We call $\Phi$ the inventory aversion parameter and we call $\mathfrak{p}$ the cutoff probability. The strategy internalises trades when the prediction $p^{\pm,\text{M}}$ is lower than the cutoff probability $\mathfrak{p}$ adjusted by the inventory of the broker $Q$ and the inventory aversion parameter $\Phi$. When either
@@
the broker is indifferent between internalising or externalising the trade in the market; this happens with probability zero.
The model variables $S$, $\eta$, and $\phi$ in (5.1) are not calibrated in our experiments. Their role here is to motivate the broker’s decision rule (5.2) in terms of the cutoff probability $\mathfrak{p}$ and the inventory aversion parameter $\Phi$.
Next, we study the case $\Phi=0$ in more detail. In C.1 we explore the case $\Phi>0$.
@@ 5.2 Internalise-externalise strategy
Motivated by the mathematical framework in Subsection 5.1, below we introduce a family of predictions of toxicity based on the cutoff probability $\mathfrak{p}\in[0,1]$. In what follows, we ignore the permanent price impact of externalising trades, as this would require the formulation of a stochastic control problem. [^7] For a series of recent related papers on broker-client stochastic control problems see Cartea and Sánchez-Betancourt (2025); Bergault and Sánchez-Betancourt (2025); Cartea et al. (2025); Donnelly and Li (2025); Wu and Jaimungal (2024). In these papers, the externalisation activity of the broker has a permanent price impact in prices. There, trades are toxic because informed traders observe the stochastic drift of the asset price and exploit that information optimally. The problem then is to characterise the equilibrium between the broker and the informed traders (the uninformed trader is usually assumed to be non-strategic). One can also allow the broker to have flexibility in the skewing mechanism, as opposed to using bid and ask quotes that are aligned to those in the lit market, see e.g., Barzykin et al. (2023, 2022). Furthermore, the broker can employ filtering techniques to identify toxic trades. For example, if skewed quotes which are “expensive” are hid by a client, this may convey information about the toxicity of the trade; see e.g., Aqsha et al. (2024) where the authors filter the trading speed of informed traders in search of the signal. Below, when deploying our strategies, the historical data do not change.
@@
We are interested in the predictive performance of the models as we vary the value of $\mathfrak{p}$. To this end, let $y_{t_{n}}\in\{0,1\}$ denote if a trade executed at time $t_{n}$ was toxic at time $t_{n}+\mathcal{G}$ ($y_{t_{n}}=1$ if toxic and $y_{t_{n}}=0$ otherwise). We employ the true positive rate and the false positive rate, which we define below.
@@
The area under an ROC curve, called AUC, is used in the machine learning literature to compare classifiers; see, e.g., Fawcett (2006). Intuitively, the AUC is a measure to quantify a classifier’s ability to distinguish between toxic and benign trades.
6 Experiments
We employ the methodology developed in the previous section with the following configuration. The NNet for PULSE is an MLP with three hidden layers, 100 units in each layer, and ReLU activation function. The number of epochs $E$ is 850, we skip the first $50$ iterations of the optimisation procedure, the subspace dimension is $d=20$, the learning rate is $\alpha=10^{-7}$, and we store gradients every $k=4$ steps. [^8] Our results are robust to higher subspace dimensions. In particular, we find that using a dimension up to $d=1000$ does not yield a statistically significant outperformance over the baseline with $d=20$. With this configuration, we estimate the $38,700$ units of the MLP during the warmup stage. For the deployment stage, PULSE updates $120$ degrees of freedom; this accounts for less than half a percent of all parameters updated during the warmup stage. From a practical perspective, the memory cost of a single step of the algorithm (during the deployment stage) is $O((L+D)^{2})$. In this paper, an update requires less than 1mb of memory because each unit is a 32bit float. Conversely, if we did not employ the subspace approach a single step would require 190gb of memory, making it infeasible to run on typical GPU devices.
We divide the dataset into a warmup dataset ${\cal D}_{\text{warmup}}$ and a deploy dataset ${\cal D}_{\text{deploy}}$. Here, ${\cal D}_{\text{warmup}}$ is from 28 June 2022 to 29 July 2022, and ${\cal D}_{\text{deploy}}$ is from 1 August 2022 to 21 October 2022.
6.1 Benchmark models
We compare the performance of four methods: PULSE, logistic regression (LogR), random forests (RF), and a recursively updated maximum-likelihood estimator of a Bernoulli-distributed random variable (MLE). The MLE benchmark reflects a common industry practice in which toxicity is treated as a client-level attribute rather than a trade-specific one. We include LogR as a widely used linear baseline that helps quantify the added value of employing non-linear models such as neural networks. Finally, we include RF, a strong tree-based nonparametric method that is competitive on tabular data (e.g., McElfresh et al., 2023).
With logistic regression, the probability that a trade is toxic is
where ${\bf w}_{0}$ is estimated maximising the log-likelihood with L-BFGS; see Liu and Nocedal (1989).
Next, RF is a bootstrap-aggregated collection of de-correlated trees. To predict if a trade is toxic one uses the average over the individual trees in the ensemble, see Section 15.1 of Hastie et al. (2001).
Further, we model the unconditional probability of a toxic trade as a Bernoulli-distributed random variable with mass function
$p(y\,|\,\pi)=\text{Bern}\Big(y\,|\,\pi\Big),\qquad y\in\{0,1\}\,.$ (6.2)
The maximum-likelihood estimator of the parameter $\pi$, given a collection $\{z_{1},\ldots,z_{N}\}$ of Bernoulli-distributed samples, where $z_{n}\in\{0,1\}$, is given by
$\pi_{\text{MLE}}=\frac{1}{N}\,\sum_{n=1}^{N}\mathbbm{1}(z_{n}=1)\,;$
The decision rule in (5.2) is directional and therefore induces two conditional problems: one for buy trades and one for sell trades. We consequently learn side-specific feature transformations and parameters. There are three implementations:
${\bf x}_{\text{bid}}\to{\cal M}_{\text{bid}}\to y,\;\;{\bf x}_{\text{ask}}\to{\cal M}_{\text{ask}}\to y,$ (6.3)
$[{\bf x}_{\text{bid}},\,{\bf x}_{\text{ask}}]\to{\cal M}\to[y_{\text{bid}},y_{\text{ask}}],$ (6.4)
$[{\bf x},\text{bid/ask}]\to{\cal M}\to y,$ (6.5)
[…]
In C.3 we study the performance when we train one model per client and use the client’s unique features.
Online evaluation. We evaluate MLE and PULSE online because both admit single-pass updates compatible with asynchronous labels. Our pipeline also supports online logistic regression; however, in our data, LogR underperforms even with batch retraining, yielding score distributions concentrated near the base rate, so online updates would not improve its behaviour. Fully online RF would require incremental trees or replay buffers which require more compute power takes longer to run. For robustness, in C.5 we re-train LogR and RF weekly on accrued data.
@@ 6.2 Model comparison
In this section, we analyse the AUC of the methods for various toxicity horizons. For each method, we compute the AUC for the sequence of trades of each day and show a density plot of such values in Figure 9; recall that the deploy window is between 1 August 2022 and 21 October 2022.
Figure 9: Left panel shows the median (line) and interquartile range (shaded) of daily AUC by toxicity horizon and model. Right panel shows the daily AUC of all models for a toxicity horizon of 30s. The dashed vertical lines correspond to the mean daily value. EUR/USD currency pair over the period 1 August 2022 to 21 October 2022.
PULSE has the highest average AUC among the four methods for all toxicity horizons. Furthermore, as the toxic horizon increases, the AUC of PULSE and RF decrease but the outperformance of PULSE over RF increases. MLE and LogR have the poorest performance. We observe that the AUC for PULSE and RF decreases as the toxicity horizon increases, reflecting that the prediction problem is more difficult over longer horizons. Empirically, the variance of profitability increases with time (see Figure 3), which reduces the predictive power of these methods. For MLE, the AUC remains close to 0.5 across horizons, highlighting that it acts as a random classifier. On the other hand, LogR shows high variability across horizons, mainly because its linear structure is not well-suited to capture the dynamics of the data, leading to poor calibration and unstable performance.
Figure 10 shows the five-day exponentially-weighted average of the AUC for each day. RF, PULSE, and MLE attain their maximum values at the beginning of the deploy period and then decay over time, while PULSE maintains a steady performance because PULSE updates its parameters with each new observation. The performance of MLE and LogR is similar. In our data, LogR does not find a useful linear boundary. Consequently, LogR’s predicted probabilities cluster near the unconditional rate and are highly correlated with MLE’s, yielding little time variation.
@@
Figure 11 shows a five-day exponentially-weighted moving average (with decay $1/3$) of the AUC for the various toxicity horizons across time.
@@
We observe that as the toxicity horizon increases, the maximum AUC decreases almost uniformly.
6.3 Trade Prediction Effectiveness and Missed Opportunities
We employ data for the EUR/USD currency pair over the period 1 August 2022 to 21 October 2022 to test the internalisation-externalisation strategy. As above, we ignore inventory aversion (i.e., $\Phi=0$).
In what follows, we define the avoided profit of an externalised trade to be the PnL of that trade had it been internalised with unwinding at the end of the toxicity horizon. Figure 12 reports the PnL ($y$-axis) of the internalised trades and the avoided profit ($x$-axis) of the externalised trades when the broker uses the internalisation-externalisation strategy (5.2) and for PULSE, MLE, LogR, and RF. The points shown for each method and toxicity horizon are the ones that maximised PnL over all possible cutoff probabilities $\mathfrak{p}\in\{0.05,0.15,0.25,\ldots,0.95\}$.
The broker starts with zero inventory at the beginning of 1 August 2022 and she crosses the spread to unwind the internalised trades at the end of the toxicity horizon. We keep track of the PnL they would have obtained over the toxicity horizon. [^9] Studying the variance of these results is more challenging and left for future research. The inventory is in euros (€) and the PnL is in dollars (\$). Each trade is for the median quantity which is €2,000 in our dataset. When a trade is toxic, the median loss to unwind the position is $\mathdollar 7\times 10^{-5}$ per euro traded and when the trade is not toxic, the median profit is $\mathdollar 8\times 10^{-5}$ per euro traded.
Figure 12: PnL and avoided profit for various toxicity horizons and for PULSE, MLE, LogR, and RF. Each point shows the highest possible PnL for a given method and toxicity horizon where the maximum is taken across cutoff probability where $\mathfrak{p}\in\{0.05,0.15,0.25,\ldots,0.95\}$.
For each toxicity horizon, the internalisation-externalisation strategy informed by PULSE attains the highest PnL and the lowest avoided profits, [^10] A negative avoided profit is an incurred loss had the trade been internalised and unwound at the toxicity horizon. see red dots joined by the dash line. These results show the added economic value that one obtains when informing trading strategy (5.2) with the predictions made by PULSE. Indeed, we show that the higher quality of predictions obtained by PULSE produce higher PnL and lower avoided profits.
Next, Figure 13 shows a histogram of $p^{\pm,\,\text{M}}$ for $\text{M}\in\{\text{{PULSE}},\text{MLE},\text{LogR},\text{RF}\}$ and toxicity horizon of 10s and Figure 14 shows how the percentage of internalised volume depends on the cutoff probability.
Figure 13: Histograms of predicted toxicity probabilities by model (pooled across clients). Toxicity horizon is ten seconds. EUR/USD currency pair over the period 1 August 2022 to 21 October 2022.
MLE produces an almost constant score (around the unconditional toxicity rate). For LogR, the concentration of probabilities below $0.5$ is due to the nonlinear structure of the data. Consequently, when the cutoff crosses these concentrated values, the internalised volume changes abruptly. In contrast, PULSE produces a broader, more dispersed (higher-entropy) distribution of probabilities, enabling smoother and more informative policy adjustments.